OpenStudy (anonymous):

7 over 5x squared -20x+20 plus 1 over 20x squared - 80

OpenStudy (anonymous):

lol, the question is not clear >_<

OpenStudy (anonymous):

7 over 5x squared minus 20x plus 20 add to 1 over 20x squared minus 80

OpenStudy (anonymous):

i agree with sstarica because 7over 5x or 7over5 then an x ?

OpenStudy (anonymous):

\[\frac{7}{5x^2} - 20x + 20 + \frac{1}{20x^2} - 80\]?

OpenStudy (anonymous):

______7_____ 5x squared -20x+20 added to 1 over 20x squared - 80

OpenStudy (anonymous):

so is sstarica right?

OpenStudy (anonymous):

sorry for the confusion Ihavent quited figured out how to work this program yet

OpenStudy (anonymous):

\[\frac{7}{5x^2-20x+20} + \frac{1}{20x-80}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright lol :)

OpenStudy (anonymous):

how do you get horizontal bars for the fractions with the maths editor

OpenStudy (anonymous):

frac{a}{b} <-- ^_^

OpenStudy (anonymous):

\[\frac{7}{5(x-4)^2} + \frac{1}{20(x-4)}\]\[\frac{7}{5(x^2-4x+4)} + \frac{1}{20(x-4)}\]

OpenStudy (anonymous):

wrong order lol

OpenStudy (anonymous):

lol, I noticed

OpenStudy (anonymous):

multiply top and bottom of second by x-4

OpenStudy (anonymous):

it's actually (x-2)^2 :)

OpenStudy (anonymous):

nvm LOL, I wrote the wrong question

OpenStudy (anonymous):

yeh something like that , im going to go to sleep soon

OpenStudy (anonymous):

no, wait, I wrote the right question >_<

OpenStudy (anonymous):

(x-2)^2 ^_^" instead of (x-4)^2

OpenStudy (anonymous):

its 1.40 in the morning where I am , I am too generous staying up this late to help ppls lol

OpenStudy (anonymous):

lol, go to bed

OpenStudy (anonymous):

\[= \frac{7(4)(x-4) + (x-2)(x-2)}{20(x-2)^2(x-4)}\] \[= \frac{28x - 112 +x^2 - 4x -4}{20(x-2)^2(x-4)}\] \[= \frac{x^2 + 24x -116}{20(x-2)^2(x-4)}\] hmm, there's something wrong

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

did you understand what I just did?

OpenStudy (anonymous):

not really i'm not very smart with algebra

OpenStudy (anonymous):

I factored, then put both fractions with the same denominator to have them in one :)

OpenStudy (anonymous):

write down the steps on paper, and you'll figure out what I did ^_^

OpenStudy (anonymous):

wait a min, there's a mistake in the last 2 steps, hold on dear ^_^

OpenStudy (anonymous):

I'm lost because under the 1 in 20x squared minues 80

OpenStudy (anonymous):

\[= \frac{28x-112+x2-4x+4}{20(x-2)^2(x-4)}\]<-- for the second line

OpenStudy (anonymous):

I factored, I pulled out 20 as a common factor and was left with (x-4) use the distributive law to check ^_^

OpenStudy (anonymous):

\[= \frac{x^2+24x-108}{20(x-2)^2(x-4)}\]

OpenStudy (anonymous):

sorry about the mess, but are you following now :)?

OpenStudy (anonymous):

i think so, thank you so much

OpenStudy (anonymous):

np ^_^, but there's still something fishy about it lol, I can't seem to factor the numerator

OpenStudy (anonymous):

does the question want you to add only? or simplify?

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

I'm sure of the following steps I just made, but looks it can't be factored

OpenStudy (anonymous):

thank you